OpenStudy (anonymous):

If f(x) = 1/9x - 2 , what is f–1(x)?

Nnesha (nnesha):

is it \[\huge\rm \frac{ 1 }{ 9x-2 } ~~~or~~\frac{ 1 }{ 9x}-2\]

Nnesha (nnesha):

??

OpenStudy (anonymous):

1/9x + 2 is one of the answer

Nnesha (nnesha):

no i'm asking which one is your question first one or 2nd ?is it \[\huge\rm \frac{ 1 }{ 9x-2 } ~~~or~~\frac{ 1 }{ 9x}-2\]

OpenStudy (anonymous):

Oh, the second one.

Nnesha (nnesha):

alright so to find inverse switch `y` and `x` variable f(x) is same as y so \[\huge\rm \color{reD}{y}=\frac{ 1 }{ 9\color{blue}{x} }-2\] inverse would be \[\huge\rm\color{blue}{x}=\frac{ 1}{ 9\color{ReD}{y}}-2\] now solve for y

Nnesha (nnesha):

do you know how to solve that equation for y?

OpenStudy (anonymous):

y = 1/9(x+ 2) ?